# Unravelling the Mechanisms of Manipulating Numbers in Language Models

### v1

@@ A.2 Generalization of probes to unseen layers
Figure 13: Probe fitted on one layer evaluated on all layers. Olmo 2 1B shows the weakest cross-layer generalization among all models. Llama models display a strong separation between input/output embedding representations and the hidden representations.
A.3 Visualizations of internal activations
@@
We visualize the first 16 PCA dimensions in Figure 14 and the maximal magnitudes of the frequencies in the Fourier transform in Figure 15. We visualize across model sizes and families.
@@
Appendix B Natural Language Dataset Description
@@
Add Sub Mul Div
Accuracy 100% 99.8% 90.4% 5.9%
P(Extracted $|$ Incorrect) - 56.8% 26.3% 94.4%
P(Not extracted $|$ Correct) 1.4% 0.03% 26.1% 5.9%
B.1 Qualitative assessment of output embeddings
((a))
((b))
Figure 16: PCA of Llama 3 1B output representations for the operations of addition (recoverable with close-to 100% accuracy) and multiplication (recoverable with appx. 50% of accuracy)
We visualize output representations of Llama 3 1B on addition and multiplication operations. For each expected result value y in the range 0-1000, we sample a random pair (x1, x2) from the same range, such that x1+x2=y (or x1*x2=y, respectively). We then predict the next token for prompts “x1 plus x2 is ” and “x1 multiplied by x2 is ” and collect the final output representations of the model before decoding. We then reduce the representations with PCA to 16 dimensions and visualize the result in Figure 16.

### v2

@@ A.3 Generalization of probes to unseen layers
Figure 12: Probe fitted on one layer evaluated on all layers, synthetic math contexts. Olmo 2 1B shows the weakest cross-layer generalization among all models. Llama models display a strong separation between input/output embedding representations and the hidden representations.
A.4 Visualizations of internal activations
@@
We visualize the first 16 PCA dimensions in Figure˜13 and the maximal magnitudes of the frequencies in the Fourier transform in Figure˜14. We visualize across model sizes and families.
@@
Add Sub Mul Div
Accuracy 100% 99.8% 90.4% 5.9%
P(Extracted $|$ Incorrect) - 56.8% 26.3% 94.4%
P(Not extracted $|$ Correct) 1.4% 0.03% 26.1% 5.9%
A.5 Qualitative assessment of output embeddings
((a))
((b))
Figure 15: PCA of Llama 3 1B output representations for the operations of addition (recoverable with close-to 100% accuracy) and multiplication (recoverable with appx. 80% of accuracy)
We visualize output representations of Llama 3 1B on addition and multiplication operations. For each expected result value $y$ in the range 0-999, we sample a random pair ($x1$, $x2$) from the same range, such that $x1+x2=y$ (or $x1\times x2=y$, respectively). We then predict the next token for prompts $x1$ plus $x2$ is ” and $x1$ multiplied by $x2$ is ” and collect the final output representations of the model before decoding. We then reduce the representations with PCA to 16 dimensions and visualize the result in Figure˜15.
A.6 Probe sparsity
This is visualized in Figure˜16, showing that intermediate layers in Llama 1B lead to probes with very few weights with values $10^{-5}$.
Nevertheless, we note that this trend is pertinent across different models to a different extent: OLMo 7B and Phi 15B learn more distributed features also across the intermediate layers — which, in turn, also leads to a better generalization of universal probes (Figure 5b), but still causing an outstanding drop in held-out accuracy compared to other layers. Finally, we note that the models’ sparsity profile does not relate to reported performances on arithmetic tasks kadlčík2025pretrainedlanguagemodelslearn and it also does not necessarily determine the accuracy of probes trained specifically for particular layer(s) (Figure˜3).
A.7 Using probes for tracking representation consistency
Figure 17: Accuracy of extracting input numbers from each component of the Transformer block using the sinusoidal probe.
We explore the origin of the high consistency of representations across layers within the internal mechanism that each transformer layer implements. In Figure 17, we visualize the accuracy of probing input representations from each of the components present in transformer layers (Llama 3.2 1B for brevity). We find the accurate numeric representation is scattered across different components in the attention mechanism, with the attention output projection largely violating the sinusoidal representation, which is then, to a large extent, reconstructed in the subsequent fully-connected block. Nevertheless, the consistency of layers’ output representation is primarily maintained by residual streams across layers.
A.8 Incremental refinement of model predictions on arithmetic tasks
((a))
((b))
((c))
((d))
We can use sinusoidal probes to estimate the absolute error per layer: the absolute difference between the predicted and target integer provides an estimate of how far off from the intended target a model is at any intermediate step of computation. Results aggregated per layer of Llama 3.2 3B are displayed in Figure˜18. Across all operations, models tend to incrementally reduce the error towards the true answer value, gradually refining across layers. These results also suggest that there are particular layers responsible for an increase of errors in the model’s internal computation.
((a))
((b))
((c))
((d))
Figure 19: Error aggregation per layer: Relative ratio of cases where each of the layers in Llama 3.2 3B breaks the correct result from the previous layer.
Figure˜19 corroborates this hypothesis — showing that particular layers ‘break’ the correct result probed from the previous layer in large proportions of all predictions. For instance, in division, the 5th layer is responsible for breaking the correct result recoverable from the previous layer in over 10% of all cases. Further analyses, presented in Table˜2, show that according to the probed representations, the models often achieve a correct result internally, even though it does not surface: this is the case of 56.8% of all surfaced (prediction) errors in subtraction, 26.3% in multiplication, and as much as 94.4% of all errors in division.
Having used our probes to identify key layers that are likely responsible for errors, we can now turn to practical interventions that they make possible. In practice, we simply remove the suspected layers from the model. This extremely naive intervention strategy is justified by the universality of representations across layers (Section˜3.3). We remove one of the three layers with the highest error aggregation, separately in multiplication and in division (having a potential for improvement of accuracy). We find that this enhances performance in four out of six cases — in one case (layer 4) in multiplication, causing a 26% error reduction (from 90.38% to 92.91% and in all cases in division, with error reduction between 27–64%. However, we note that the absolute number of correct predictions probed in the lower layers is negligible, which hinders the general applicability of this methodology.
Appendix B Experimental details
The overview provided here is intended for informative purposes. For precise replication, we refer the reader to the companion codebase, which contains the exact hyperparameters for each experiment, available at https://github.com/prompteus/numllama.
In our probe architecture param-sin, we initialize amplitude parameters $\mathbf{w}_{a}$ with a constant of $0.01$, phase-shift parameters $\mathbf{w}_{\phi}$ randomly from uniform distribution between 0 and $2\pi$, and frequency parameters $\mathbf{w}_{f}$ with a vector of values regularly-spaced between 0 and $\frac{\pi}{4}$, and the parameters of linear projection $\mathbf{W}_{\mathrm{in}}$ randomly from uniform distribution between $\sqrt{1/n}$ and $\sqrt{1/n}$, where $d$ is the input dimension (number of model features).
We find that L1-regularization of the amplitude parameters $\mathbf{w}_{a}$ improves generalization to held-out tokens by encouraging the probe to maintain small effective latent dimensionality even with a large actual dimension $d$. Similarly, we found similar benefits from regularization of the linear projection $\mathbf{W}_{\mathrm{in}}$, although L2-regularization appears more robust. We find that an analogous configuration is also effective for the probe of kadlčík2025pretrainedlanguagemodelslearn, with L2-regularization of $\mathbf{W}_{\mathrm{in}}$ and L1-regularization of $\mathbf{W}_{\mathrm{out}}$ for inducing simplicity (sparsity) in the sinusoidal basis.
The probes are fitted on a train set, with a checkpoint selected on the validation set, and then evaluated on the held-out test set. In most experiments, we use the Adam optimizer with a learning rate of $10^{-4}$ and apply regularization with a coefficient of either $10^{-3}$ or $10^{-2}$.
In multi-token decoding (Figure˜6), we use a learning rate of $5\cdot 10^{-4}$, 10,000 training steps, and early-stop training.
When fitting probes on natural language contexts (Figure˜4, ‘Nat. lang. $\rightarrow$ Nat. lang.’) and for cross-layer transfer (Figure˜12), we use a hidden dimension of $d=500$.
In our steering experiments (Figure˜9), for $f_{\sin}$, we optimize embeddings with respect to the probe using an SGD optimizer with a learning rate of $0.1$, optimize the embeddings for 200,000 steps, and decrease the learning rate by a factor of 100 over the first 100,000 steps. As for $f_{\mathrm{param}\sin}$, we set regularization coefficient to $10^{-4}$ and use a hidden dimension of $d=2000$.
@@ Appendix B Experimental details
We provide an overview of the datasets we use for natural language context experiments in Table˜3. In Section˜4.1, we use BBC data as processed by Li et al. (2024), which we retrieved from RealTimeData/bbc_news_alltime.
In all cases, we trust the original creators of the datasets for handling private and sensitive information according to standard. As we do not redistribute the data, we take no specific steps towards verifying this point.
Appendix C Use of LLMs in writing and coding
LLM-based technology was used to polish writing and jump-start coding. Authors take full responsibility for the contents of this paper.